OpenStudy (anonymous):

under the QUADRATIC EQUATION(Roots) If a triangle has an area of 56 sq. inches, with base of 9 inches longer that its altitude, find the length of its base.

OpenStudy (anonymous):

area of triangle = 0.5 * base * length let length of base = x then height = x-9 so 56 = 0.5 * x(x-9) can u solve this?

OpenStudy (anonymous):

Let h = altitude of triangle h+9 = base of triangle \[h(h+9) = 56\]\[h^2+9h-56=0\]\[h=\frac{-9\pm \sqrt{36+4(56)}}{2}\]\[h=\frac{-9\pm \sqrt{260}}{2}\] But no negative values are allowed in geometry. Therefore: \[h=\frac{-9+ \sqrt{260}}{2}\]

OpenStudy (anonymous):

(1/2)*b*h =56 b=h+9 =>(1/2)*(h+9)*h=56 => h^2 + 9h -112 =0 =>h=7 , -16 as h is positive h=7 => b = h+9 =16

OpenStudy (anonymous):

According to Wolfram, this is the answer: 3.56225 http://www.wolframalpha.com/input/?i=%5B-9%2Bsqrt%28260%29%5D%2F2

OpenStudy (anonymous):

its a quadratic equation rearrange 0.5x^2 - 4.5x - 56 = 0 multiply by 2 to remove decimals x^2 - 9x - 112 = 0 solve by factoring - if u cant use the formula

OpenStudy (anonymous):

solution is 16 or -7 length cant be nagative so x = 16 length of base = 16 ins

OpenStudy (anonymous):

Whoops forgot about the (1/2)*bh formula.

OpenStudy (anonymous):

you made a slight error in ur calculation jyqft on third line - should be 81 not 36

OpenStudy (anonymous):

i used the formula also but i've just seen that it will factor: x^2 - 9x - 112 = (x-16)(x+7)

OpenStudy (anonymous):

ok with that kyra?

OpenStudy (anonymous):

how is manila these days?